OpenStudy (anonymous):

find the derivative step by step of y = 2^x

OpenStudy (anonymous):

y = x^(2x) ==> ln(y) = ln[x^(2x)] ==> ln(y) = 2x*ln(x). Then by logarithmic differentiation and the Product Rule: 1/y * dy/dx = 2ln(x) + 2x(1/x) ==> 1/y * dy/dx = 2[ln(x) + 1] ==> dy/dx = 2y * [ln(x) + 1] ==> dy/dx = 2 * x^(2x) * [ln(x) + 1].

OpenStudy (cwrw238):

lake logs: ln y = ln 2^x now differentiate both sides and simplify

OpenStudy (cwrw238):

blast - is yore first line correc?

OpenStudy (anonymous):

ny = 2x*lnx 1/y dy/dx = 2lnx + 2 dy/dx = 2y(lnx + 1) dy/dx = 2* x^(2x) * (lnx + 1)

OpenStudy (anonymous):

Thanx for correcting me

OpenStudy (cwrw238):

THATS OK